The cohesive energy of a sodium chloride crystal is 6eV per ion pair. The ionization energy of Na is 5eV and the electron affinity of Cl is 3.75eV. Neglecting the energy of repulsion, estimate the interionic spacing. (The electron affinity of Cl is the binding energy of the last electron of Cl?).
Answer #1

media%2F478%2F47853878-2eae-4e65-8baf-38

Adding all the three, E + 3.75 - 5 = 6 ; E = 7.25 eV


Let the inter ionic spacing be d. and (since repulsion is neglected)


This potential energy = k*q1*q2/d ; where q1 and q2 are charges on the two ions , each equal to 1e in this case.


7.25 * 1.6*10^-19 = k*e*e/d

d = 1/[7.25/(k*e)] = 1.986 * 10^-10 m = 2 Angstrom